OpenStudy (anonymous):

Which shows 532 - 472 being evaluated using the difference of squares method? 532 - 472 = 2809 - 2209 = 600 532 - 472 = (53 + 47)(53 - 47) = (100)(6) = 600 532 - 472 = (53 - 47)2 = 62 = 36 532 - 472 = (2809 + 2209)(2809 - 2209) = 3,010,800

OpenStudy (anonymous):

Which shows 53^2 - 47^2 being evaluated using the difference of squares method? 53^2 - 47^2 = 2809 - 2209 = 600 53^2 - 47^2 = (53 + 47)(53 - 47) = (100)(6) = 600 53^2 - 47^2 = (53 - 47)^2 = 6^2 = 36 53^2 - 47^2 = (2809 + 2209)(2809 - 2209) = 3,010,800 sorry posted and didnt relize it didnt put the powers in correctly

OpenStudy (whpalmer4):

Difference of squares: \[a^2-b^2 = (a+b)(a-b)\] You have \(a = 53\) and \(b=47\)